# ADELIE: Aligning Large Language Models on Information Extraction

### v1

@@ Abstract
Large language models (LLMs) usually fall short on information extraction (IE) tasks and struggle to follow the complex instructions of IE tasks. This primarily arises from LLMs not being aligned with humans, as mainstream alignment datasets typically do not include IE data. In this paper, we introduce ADELIE (Aligning large language moDELs on Information Extraction), an aligned LLM that effectively solves various IE tasks, including closed IE, open IE, and on-demand IE. We first collect and construct a high-quality alignment corpus IEInstruct for IE. Then we train ADELIESFT using instruction tuning on IEInstruct. We further train ADELIESFT with direct preference optimization (DPO) objective, resulting in ADELIEDPO. Extensive experiments on various held-out IE datasets demonstrate that our models (ADELIESFT and ADELIEDPO) achieve state-of-the-art (SoTA) performance among open-source models. We further explore the general capabilities of ADELIE, and experimental results reveal that their general capabilities do not exhibit a noticeable decline. We will release the code, data, and models to facilitate further research. [^1] https://github.com/THU-KEG/ADELIE
@@
Yunjia Qithanks: Equal contribution., Hao Peng , Xiaozhi Wang, Bin Xu, Lei Hou, Juanzi Li Tsinghua University, Beijing, China {qyj23, peng-h21}@mails.tsinghua.edu.cn
@@ 5.2 Experimental Results
The results on held-out closed IE datasets are shown in Table 1. We can observe that: (1) ADELIESFT performs significantly better than the original LLAMA 2 and surpasses all IE LLMs and GPT-3.5, even on par with GPT-4. This indicates that alignment on our rich data IEInstruct is effective. (2) DPO further enhances performance. ADELIEDPO performs consistently better than ADELIESFT across most datasets. This suggests that for extractive tasks with ground truth answers, further alignment using DPO can also self-improve model performance. However, the improvement of DPO is generally modest, possibly due to not using additional human-annotated preference pairs. We leave using human-annotated preference pairs for training DPO as future work. (3) Incorporating in-context demonstrations during the alignment process is necessary. Previous work only focuses on zero-shot capabilities and overlooks few-shot capabilities of LLMs, resulting in no significant improvement or even a decline when providing few-shot demonstrations, e.g., a $4.3\%$ decline in F1 score for GoLLIE. In contrast, ADELIESFT ’s few-shot performance is much better than its zero-shot performance, which suggests that ADELIESFT possesses few-shot in-context learning capabilities for closed IE tasks. It demonstrates the effectiveness of including in-context demonstrations in the alignment process.

### v2

@@ Abstract
Large language models (LLMs) usually fall short on information extraction (IE) tasks and struggle to follow the complex instructions of IE tasks. This primarily arises from LLMs not being aligned with humans, as mainstream alignment datasets typically do not include IE data. In this paper, we introduce ADELIE (Aligning large language moDELs on Information Extraction), an aligned LLM that effectively solves various IE tasks, including closed IE, open IE, and on-demand IE. We first collect and construct a high-quality alignment corpus IEInstruct for IE. Then we train ADELIESFT using instruction tuning on IEInstruct. We further train ADELIESFT with direct preference optimization (DPO) objective, resulting in ADELIEDPO. Extensive experiments on various held-out IE datasets demonstrate that our models (ADELIESFT and ADELIEDPO) achieve state-of-the-art (SoTA) performance among open-source models. We further explore the general capabilities of ADELIE, and experimental results reveal that their general capabilities do not exhibit a noticeable decline. We have released the code, data, and models to facilitate further research. [^1] https://github.com/THU-KEG/ADELIE
@@
Yunjia Qithanks: Equal contribution., Hao Peng , Xiaozhi Wang, Bin Xuthanks: Corresponding author: xubin@tsinghua.edu.cn, Lei Hou, Juanzi Li Department of Computer Science and Technology, BNRist, Tsinghua University {qyj23, peng-h24}@mails.tsinghua.edu.cn
@@ 5.2 Experimental Results
The results on held-out closed IE datasets are shown in Table 1. We can observe that: (1) ADELIESFT performs significantly better than the original LLAMA 2 and surpasses all IE LLMs and GPT-3.5, on par with GPT-4. Compared to InstructUIE and GoLLIE, which adopt more advanced base LLMs (FLAN-T5 11B and Code LLAMA 7B ) in IE tasks (Peng et al., 2023a) and more SFT data ($144$k and $165$k IE instances), ADELIESFT achieves better results using only $83$k SFT data with LLAMA 2 7B. This indicates that our data construction method is effective and IEInstruct is of high quality. (2) DPO further enhances performance. ADELIEDPO performs consistently better than ADELIESFT across most datasets. This suggests that for extractive tasks with ground truth answers, further alignment using DPO can also self-improve model performance. However, the improvement of DPO is generally modest, possibly due to not using additional human-annotated preference pairs. We leave using human-annotated preference pairs for training DPO as future work. (3) Incorporating in-context demonstrations during the alignment process is necessary. Previous work only focuses on zero-shot capabilities and overlooks few-shot capabilities of LLMs, resulting in no significant improvement or even a decline when providing few-shot demonstrations, e.g., a $4.3\%$ decline in F1 score for GoLLIE. In contrast, ADELIESFT ’s few-shot performance is much better than its zero-shot performance, which suggests that ADELIESFT possesses few-shot in-context learning capabilities for closed IE tasks. It demonstrates the effectiveness of including in-context demonstrations in the alignment process.
@@ 7 Conclusion
Acknowledgements
We thank all the anonymous reviewers and meta reviewers for their valuable comments. This work is supported by the National Natural Science Foundation of China (No. 62277033), Beijing Natural Science Foundation (L243006), a grant from the Institute for Guo Qiang, Tsinghua University (2019GQB0003) and the project from Tsinghua-SPD Bank Joint-Lab. Thanks the support from National Engineering Laboratory for Cyberlearning and Intelligent Technology, and Beijing Key Lab of Networked Multimedia.